# Towards Universal Dense Blocking for Entity Resolution

### v1

@@ Abstract.
Blocking is a critical step in entity resolution, and the emergence of neural network-based representation models has led to the development of dense blocking as a promising approach for exploring deep semantics in blocking. However, previous advanced self-supervised dense blocking approaches require domain-specific training on the target domain, which limits the benefits and rapid adaptation of these methods. To address this issue, we propose UBlocker, a dense blocker that is pre-trained on a domain-independent, easily-obtainable tabular corpus using self-supervised contrastive learning. By conducting domain-independent pre-training, UBlocker can be adapted to various downstream blocking scenarios without requiring domain-specific fine-tuning. To evaluate the universality of our entity blocker, we also construct a new benchmark covering a wide range of blocking tasks from multiple domains and scenarios. Our experiments show that the proposed UBlocker, without any domain-specific learning, significantly outperforms previous self- and unsupervised dense blocking methods and is comparable and complementary to the state-of-the-art sparse blocking methods.
@@ 1. Introduction
Unfortunately, recent studies (Papadakis et al., 2022; Zeakis et al., 2023) have shown that, in contrast to supervised dense blocking methods (Zhang et al., 2020; Brinkmann et al., 2023), self-supervised dense blocking methods (Thirumuruganathan et al., 2021; Wang et al., 2022) do not offer substantial benefits and may be inferior to the direct use of pre-trained word or sentence embeddings. As illustrated in Figure 1(a), we argue that this is because these methods learn domain-specific models, which means that for each new blocking task, a unique representation model needs to be trained using the available records. Dense blocking methods in this paradigm demand considerable time and computational resources for each blocking. Learning on domain-specific data can also lead to overfitting or collapse, making the outcomes do not worth the effort. However, the potential of self-supervised dense blocking learning from extensive data remains largely unexplored.
@@
In this paper, we aim to extend the methodology of dense blocking from the one-model-per-dataset to a more generalized one-model-all-datasets paradigm, as shown in Figure 1. Motivated by recent advances in large language models (Raffel et al., 2020; Brown et al., 2020), we develop a universal dense blocker, UBlocker, by exploiting the power of cross-domain self-supervised learning to pre-train on large amounts of tabular data. Figure 2 shows the overall pre-training procedure of UBlocker. UBlocker is built upon a Transformer-based architecture and trained with contrastive learning on GitTables (Hulsebos et al., 2022), a large-scale relational tabular corpus. Specifically, we filter and process the tables from this corpus into records, generate positive and negative pairs using data paraphrasing techniques at different levels, and obtain record representations by passing them through the encoder. The contrastive learning algorithm refines the model’s capability to represent records. Based on the aforementioned process, UBlocker, pre-trained on a heterogeneous, multi-domain, and multi-topic corpus of 1 million tables, acquires the capability to universally block entity collections across diverse domains and scenarios without further domain-specific learning.
@@
To ascertain the effectiveness and universality of UBlocker, we gathered and constructed a new benchmark for universal blocking evaluation. The benchmark comprises 19 datasets spread across 9 domains and 4 scenarios, which enable a thorough evaluation of the performance of UBlocker and other blocking techniques. Unlike previous studies that evaluate methods on partially labeled datasets and typically with identical schema records, this new benchmark emphasizes the ablility of a blocking technique to transfer seamlessly across domains and scenarios without requiring extra resources. As a result, the constructed benchmark evaluates entity blockers based on both effectiveness and scalability, considers both homogeneous and heterogeneous entity record collections (Papadakis et al., 2021a), and includes datasets of different scales from various domains.
@@
Thorough experiments have been carried out on the constructed benchmark. The experimental results highlight the effectiveness and universality of UBlocker, verifying the feasibility and benefits of universal dense blocking. Specifically, experiments show that UBlocker, without any additional domain-specific training, outperforms state-of-the-art (SOTA) self- and unsupervised dense blocking systems by almost 10% in mean average precision. Furthermore, UBlocker is comparable to SOTA sparse blocking methods and is more effective in some complex and dirty scenarios where duplicate records are semantically similar. In addition, UBlocker demonstrated similar and even better efficiency when blocking large volumes of records. Overall, these experimental results provide strong empirical support for the effectiveness, universality, and scalability of UBlocker.
@@
  • We provide a new insight into the limitations of existing self-supervised dense blocking methods that follow the one-model-per-dataset paradigm, and present our vision for one-model-all-datasets universal dense blocking.
  • We propose a unified contrastive pre-training framework for developing the universal dense blocker, UBlocker. We design several optimizations to better represent structural records.
  • We construct a new benchmark for universal blocking evaluation, which covers a wide range of blocking tasks from multiple domains and scenarios. This also lays the foundation for future reseach on automatic parameter search or good attribute identification for blocking (Papadakis et al., 2022; Paulsen et al., 2023).
  • We conduct thorough experiments  [^1] We will release the code and the benchmark once the paper is accepted. to demonstrate the effectiveness and universality of UBlocker by comparing it with the state-of-the-art dense and sparse blocking systems.
@@ 2. Preliminaries and Background
In this section, we first introduce the problem definition and dense blocking. Then we discuss the limitations of benchmarks for blocking.
@@ 2.1. Problem Definition
Entity resolution involves identifying records that refer to the same real-world entity, often called matches or duplicates. To avoid the quadratic number of comparisons with complicated entity matching algorithms, the blocking step is introduced in ER systems, which uses basic features to quickly scan all record pairs to achieve a reasonable trade-off between the number of comparisons and the loss of performance. Formally, the blocking step is defined in Definition 2.1.
@@ 2.3. Benchmarks for Blocking
3. UBlocker: A Universal Dense Blocker for Entity Blocking
In this section, we propose a contrastive pre-training framework to develop UBlocker, a universal dense blocker capable of blocking records across domains without domain-specific learning.
@@ 3.1. Overview
Figure 2 shows the overall pre-training procedure of UBlocker, which is a deep Transformer-based neural network trained on a large-scale relational tabular corpus using the contrastive learning algorithm. Specifically, UBlocker starts by performing column detection and data cleaning to transform tables into sets of records. Subsequently, each record is transformed into a sequence of embeddings, and a Transformer-based encoder is employed to generate a dense semantic vector. To learn an effective record encoder, we leverage data paraphrasing techniques to generate positive and negative record pairs and contrastive learning is conducted by pulling positive pairs closer while pushing negative pairs further apart.
In the following, we will explain the process of acquiring the pre-training data from the tabular corpus. We will then describe the model architecture of UBlocker and demonstrate the approach by which UBlocker is trained using self-supervised contrastive learning on tabular records.
Figure 2. Overview of UBlocker pre-training procedure.
@@ 3.2. Pre-training Data Conditioning
The divergence between unstructured natural language and structured records (i.e., tuples of attributes and values) makes the direct application of pre-trained natural language models for table-related tasks suboptimal (Dong et al., 2022a). To fill this gap, in this paper, we propose to address this problem by pre-training UBlocker on large-scale relational tables to align the unstructured natural language models with structured and semi-structured records.
@@
After cleaning the corpus, we transform all tables into sets of entity records. Each row in a table is regarded as a record. These sets of records are then fed into the UBlocker to learn a universal entity blocker, as illustrated below.
@@ 3.4. Data Paraphrasing for Positive Pair Generation
From another perspective, duplicate detection can be seen as a weak supervised learning based on sparse similarity measures. Since the blocking phase cannot afford such computationally intensive measures, models can be trained in this way to generate similar representations for similar records across different measures.
@@ 3.5. Model Training
In the following, we will describe how we train UBlocker with the aforementioned techniques. Formally, given a batch of $N$ records donated as $E={\{e_{i}\}}_{1\leq i\leq N}$, we generate positive samples for each of them by applying literal paraphrasing, resulting in $E^{\prime}={\{e_{i}^{\prime}\}}_{1\leq i\leq N}$. We then apply record embedding to these $2N$ instances and feed them into the encoder with feature paraphrasing, obtaining their representations $\boldsymbol{Z}={\{\boldsymbol{z}_{i}\}}_{1\leq i\leq N}$ and $\boldsymbol{Z}^{\prime}={\{\boldsymbol{z}_{i}^{\prime}\}}_{1\leq i\leq N}$. For the constructed pairs from these $2N$ records, we consider the pairs generated through data paraphrasing and duplicate detection as positive pairs, while the remaining pairs in the batch are regarded as negative pairs. Finally, since each instance may have multiple positive samples, we adapt Circle Loss (Sun et al., 2020) as our loss function. Suppose for each sample $e$ and its representation $\boldsymbol{z}$ we have $J$ positives and $K$ negatives, where $J+K=N$. We denote the similarity of their paraphrased version representations to $\boldsymbol{z}$ are ${\{s_{p}^{j}\}}_{1\leq j\leq J}$ and ${\{s_{n}^{k}\}}_{1\leq k\leq K}$ respectively. The loss function is described as:
@@ 4. Benchmarking Universal Blocking
This section presents a new benchmark and evaluation to overcome the limitations of benchmarks for blocking, as discussed in Section 2.3. This aims to assess the universality of blocking methods.
@@ 4.1. Overview of Proposed Benchmark
By incorporating the diversity of these factors, our benchmark provides a comprehensive evaluation of blocking methods. This also lays the foundation for future reseach on automatic parameter search or good attribute identification for blocking (Papadakis et al., 2022; Paulsen et al., 2023).
Table 2 provides an overview of the datasets included in our benchmark, which comprises a total of 17 datasets from 9 different domains. These datasets include 6 from one source, 10 from two sources and 1 from three sources. For the datasets sourced from two sources, we include 5 homogeneous and 5 heterogeneous datasets. The number of attributes per entity record ranges from 1 to 31, with an average number of duplicates per record ranging from 0.02 to 1. Different “notebook” and “altosight” datasets originate from different years of SIGMOD Programming Contest. After preprocessing and labeling, most datasets are found to be homogeneous, especially those from DeepMatcher (Mudgal et al., 2018). For the heterogeneous datasets, we include both artificially heterogeneous datasets, constructed by renaming and merging attributes as (Wang et al., 2021b), and naturally heterogeneous datasets where raw data was available. For example, we constructed “fodors-zagats_heter” and “abt-buy_heter” by renaming and merging attributes. In contrast, “walmart-amazon_heter” comprises the original heterogeneous records with more than 15 attributes and labels from “walmart-amazon_homo”. The multi-source “movies” dataset is a merge of three datasets (IMDB-TMDB, IMDB-TVDB, and TMDB-TVDB) obtained from JedAIToolkit (Papadakis et al., 2020). Details of the datasets and processing can be found in the supplementary material.
@@ 5. Experiments
DeepBlocker Sudowoodo STransformer UBlocker
@@
In the following section, we conduct thorough experiments to determine the effectiveness and universality of UBlocker, as well as to identify its strengths and weaknesses in comparison to sparse blocking methods. We first evaluate the performance of UBlocker against both domain-specific and unified sentence-based dense blocking methods. Then we compare UBlocker with sparse blocking methods. Finally, we conduct ablation studies.
@@ 5.1. Experimental Setup
We implemented UBlocker using Python 3.10, Pytorch 1.13, and the Transformers (Wolf et al., 2020) library. We chose to use the all-mpnet-base-v2 from STransformer (Reimers and Gurevych, 2019) as the initial pre-training weight and set the maximum sequence length to 256. We set the probability of applying literal paraphrasing actions to 0.01, the dropout probability for feature paraphrasing to 0.15, and the similarity threshold $s$ for duplicate detection to 0.85. The scaling factor $\gamma$ and margin $m$ of the Circle Loss are the default 80 and 0.4, respectively. We used the AdamW optimizer with a learning rate of 1e-5, a linear learning rate schedule without warm-up steps and a batch size of 128 for pre-training. The total number of pre-training steps is 10,000.
@@ 5.2. Comparison with Dense Blocking Methods
UBlocker vs Domain-Specific Methods.
To assess the effectiveness and universality of UBlocker, we first compare UBlocker with SOTA domain-specific self-supervised dense blocking methods. Experimental results presented in Table 3 indicate that UBlocker outperforms previous domain-specific dense blocking methods by almost 10% mAP on average. In contrast to the lack of a clear winner between DeepBlocker and Sudowoodo, UBlocker consistently delivers the best performance among them on almost all datasets, futher proving its effectiveness by reducing the number of nearest neighbors required. These results indicate that UBlocker not only addresses the problem of limited downstream data, but also eliminates the need for domain-specific model training. Pre-training allows deep models to reach their full potential and learn a more general representation of the structured records. In conclusion, our experimental results provide compelling evidence for the advantage of self-supervised dense blocking learning on large-scale open-domain data.
Finding 1. UBlocker outperforms SOTA domain-specific dense blockers, demonstrating the feasibility of developing universal dense blockers via a unified domain-independent pre-training.
@@
Blocking Workflows Sparkly3 UBlocker UBlocker + Sparkly3
@@
UBlocker vs Unified Sentence-Based Models.
To explore the necessity of pre-training on structured data, we compare UBlocker with the sentence embedding language model. The performance of STransformer, which served as our initialized model for pre-training, is detailed in Table 3. Our experimental results show that UBlocker outperforms STransformer by over 10% mAP on average. STransformer performs relatively poorly on some less challenging but more structured datasets, such as “fodors-zagats_homo”, “abt-buy_homo”, and “walmart-amazon_homo”. This highlights the barriers to applying language models directly to blocking due to the divergences between unstructured natural language and structured records. Overall, our results emphasize the necessity of training the universal blocker on structured and semi-structured records.
Finding 2. UBlocker outperforms sentence-based models, supporting the need to align unstructured language models with structured and semi-structured records by pre-training.
Domain-Specific Fine-tuning for UBlocker.
To investigate the potential benefits of domain-specific fine-tuning for UBlocker, we further fine-tune UBlocker in the one-model-per-dataset paradigm. The mAP changes of the fine-tuned models are also presented in Table 3. Comparing the experimental results with and without fine-tuning, we observe that UBlocker derives limited benefit from fine-tuning except for “imdb-dbpedia”. This finding suggests that the domain-independent pre-training of UBlocker has resulted in a universal blocking capability that is sufficient to address domain-specific blocking challenges without the need for further fine-tuning. Consequently, this finding reaffirms the effectiveness of UBlocker in solving blocking problems across multiple domains.
Finding 3. UBlocker benefits little from domain-specific self-supervised fine-tuning, indicating that UBlocker can learn sufficient knowledge from domain-independent pre-training.

### v2

@@ Abstract.
Blocking is a critical step in entity resolution, and the emergence of neural network-based representation models has led to the development of dense blocking as a promising approach for exploring deep semantics in blocking. However, previous advanced self-supervised dense blocking approaches require domain-specific training on the target domain, which limits the benefits and rapid adaptation of these methods. To address this issue, we propose UniBlocker, a dense blocker that is pre-trained on a domain-independent, easily-obtainable tabular corpus using self-supervised contrastive learning. By conducting domain-independent pre-training, UniBlocker can be adapted to various downstream blocking scenarios without requiring domain-specific fine-tuning. To evaluate the universality of our entity blocker, we also construct a new benchmark covering a wide range of blocking tasks from multiple domains and scenarios. Our experiments show that the proposed UniBlocker, without any domain-specific learning, significantly outperforms previous self- and unsupervised dense blocking methods and is comparable and complementary to the state-of-the-art sparse blocking methods.
@@ 1. Introduction
Unfortunately, recent studies (Papadakis et al., 2022; Zeakis et al., 2023) have shown that, in contrast to supervised dense blocking methods (Zhang et al., 2020; Brinkmann et al., 2023), self-supervised dense blocking methods (Thirumuruganathan et al., 2021; Wang et al., 2022) do not offer substantial benefits and may be inferior to the direct use of pre-trained word or sentence embeddings. As illustrated in Figure 1(a), we argue that this is because these methods learn domain-specific models, which means that for each new blocking task, a unique representation model needs to be trained using the available records. Dense blocking methods in this paradigm demand considerable time and computational resources for each blocking. Learning on domain-specific data can also lead to overfitting or collapse, making the outcomes not worth the effort. However, the potential of self-supervised dense blocking learning from extensive data remains largely unexplored.
@@
In this paper, we aim to extend the methodology of dense blocking from the one-model-per-dataset to a more generalized one-model-all-datasets paradigm, as shown in Figure 1. Motivated by recent advances in large language models (Raffel et al., 2020; Brown et al., 2020), we develop a universal dense blocker, UniBlocker, by exploiting the power of cross-domain self-supervised learning to pre-train on large amounts of tabular data. Figure 2 shows the overall pre-training procedure of UniBlocker. UniBlocker is built upon a Transformer-based architecture and trained with contrastive learning on GitTables (Hulsebos et al., 2022), a large-scale relational tabular corpus. Specifically, we filter and process the tables from this corpus into records, generate positive and negative pairs using data paraphrasing techniques at different levels, and obtain record representations by passing them through the encoder. The contrastive learning algorithm refines the model’s capability to represent records. Based on the aforementioned process, UniBlocker, pre-trained on a heterogeneous, multi-domain, and multi-topic corpus of 1 million tables, acquires the capability to universally block entity collections across diverse domains and scenarios without further domain-specific learning.
@@
To ascertain the effectiveness and universality of UniBlocker, we gathered and constructed a new benchmark for universal blocking evaluation. The benchmark comprises 19 datasets spread across 9 domains and 4 scenarios, which enable a thorough evaluation of the performance of UniBlocker and other blocking techniques. Unlike previous studies that evaluate methods on partially labeled datasets and typically with identical schema records, this new benchmark emphasizes the ability of a blocking technique to transfer seamlessly across domains and scenarios without requiring extra resources. As a result, the constructed benchmark evaluates entity blockers based on both effectiveness and scalability, considers both homogeneous and heterogeneous entity record collections (Papadakis et al., 2021a), and includes datasets of different scales from various domains.
@@
Thorough experiments have been carried out on the constructed benchmark. The experimental results highlight the effectiveness and universality of UniBlocker, verifying the feasibility and benefits of universal dense blocking. Specifically, experiments show that UniBlocker, without any additional domain-specific training, outperforms state-of-the-art (SOTA) self- and unsupervised dense blocking systems by almost 10% in mean average precision. Furthermore, UniBlocker is comparable to SOTA sparse blocking methods and is more effective in some complex and dirty scenarios where duplicate records are semantically similar. In addition, UniBlocker demonstrated similar and even better efficiency when blocking large volumes of records. Overall, these experimental results provide strong empirical support for the effectiveness, universality, and scalability of UniBlocker.
@@
  • We provide new insight into the limitations of existing self-supervised dense blocking methods that follow the one-model-per-dataset paradigm, and present our vision for one-model-all-datasets universal dense blocking.
  • We propose a unified contrastive pre-training framework for developing the universal dense blocker, UniBlocker. We design several optimizations to represent structural records better.
  • We construct a new benchmark for universal blocking evaluation, which covers a wide range of blocking tasks from multiple domains and scenarios. This also lays the foundation for future research on automatic parameter search or good attribute identification for blocking (Papadakis et al., 2022; Paulsen et al., 2023).
  • We conduct thorough experiments to demonstrate the effectiveness and universality of UniBlocker by comparing it with the state-of-the-art dense and sparse blocking systems.
@@ 2. Preliminaries and Background
In this section, we first introduce the problem definition and dense blocking. Then we discuss the limitations of blocking benchmarks.
@@ 2.1. Problem Definition
ER involves identifying records that refer to the same real-world entity, often called matches or duplicates. To avoid the quadratic number of comparisons with complicated entity matching algorithms, the blocking step is introduced in ER systems, which uses basic features to quickly scan all record pairs to achieve a reasonable trade-off between the number of comparisons and the loss of performance. Formally, the blocking step is defined in Definition 2.1.
@@ 2.3. Benchmarks for Blocking
3. UniBlocker: A Universal Dense Blocker for Entity Blocking
In this section, we propose a contrastive pre-training framework to develop UniBlocker, a universal dense blocker capable of blocking records across domains without domain-specific learning.
@@ 3.1. Overview
Figure 2 shows the overall pre-training procedure of UniBlocker, which is a deep Transformer-based neural network trained on a large-scale relational tabular corpus using the contrastive learning algorithm. Specifically, UniBlocker starts by performing column detection and data cleaning to transform tables into sets of records. Subsequently, each record is transformed into a sequence of embeddings, and a Transformer-based encoder is employed to generate a dense semantic vector. To learn an effective record encoder, we leverage data paraphrasing techniques to generate positive and negative record pairs and contrastive learning is conducted by pulling positive pairs closer while pushing negative pairs further apart.
In the following, we will explain the process of acquiring the pre-training data from the tabular corpus. We will then describe the model architecture of UniBlocker and demonstrate the approach by which UniBlocker is trained using self-supervised contrastive learning on tabular records.
Figure 2. Overview of UniBlocker pre-training procedure.
@@ 3.2. Pre-training Data Conditioning
The divergence between unstructured natural language and structured records (i.e., tuples of attributes and values) makes the direct application of pre-trained natural language models for table-related tasks suboptimal (Dong et al., 2022a). To fill this gap, in this paper, we propose to address this problem by pre-training UniBlocker on large-scale relational tables to align the unstructured natural language models with structured and semi-structured records.
@@
After cleaning the corpus, we transform all tables into sets of entity records. Each row in a table is regarded as a record. These sets of records are then fed into the UniBlocker to learn a universal entity blocker, as illustrated below.
@@ 3.4. Data Paraphrasing for Positive Pair Generation
From another perspective, duplicate detection can be seen as weak supervised learning based on sparse similarity measures. Since the blocking phase cannot afford such computationally intensive measures, models can be trained in this way to generate similar representations for similar records across different measures.
@@ 3.5. Model Training
In the following, we will describe how we train UniBlocker with the aforementioned techniques. Formally, given a batch of $N$ records donated as $E={\{e_{i}\}}_{1\leq i\leq N}$, we generate positive samples for each of them by applying literal paraphrasing, resulting in $E^{\prime}={\{e_{i}^{\prime}\}}_{1\leq i\leq N}$. We then apply record embedding to these $2N$ instances and feed them into the encoder with feature paraphrasing, obtaining their representations $\boldsymbol{Z}={\{\boldsymbol{z}_{i}\}}_{1\leq i\leq N}$ and $\boldsymbol{Z}^{\prime}={\{\boldsymbol{z}_{i}^{\prime}\}}_{1\leq i\leq N}$. For the constructed pairs from these $2N$ records, we consider the pairs generated through data paraphrasing and duplicate detection as positive pairs, while the remaining pairs in the batch are regarded as negative pairs. Finally, since each instance may have multiple positive samples, we adapt Circle Loss (Sun et al., 2020) as our loss function. Suppose for each sample $e$ and its representation $\boldsymbol{z}$ we have $J$ positives and $K$ negatives, where $J+K=N$. We denote the similarity of their paraphrased version representations to $\boldsymbol{z}$ are ${\{s_{p}^{j}\}}_{1\leq j\leq J}$ and ${\{s_{n}^{k}\}}_{1\leq k\leq K}$ respectively. The loss function is described as:
@@ 4. Benchmarking Universal Blocking
This section presents a new benchmark and evaluation to overcome the limitations of blocking benchmarks, as discussed in Section 2.3. This aims to assess the universality of blocking methods.
@@ 4.1. Overview of Proposed Benchmark
By incorporating the diversity of these factors, our benchmark provides a comprehensive evaluation of blocking methods. This also lays the foundation for future research on automatic parameter search or good attribute identification for blocking (Papadakis et al., 2022; Paulsen et al., 2023).
Table 2 provides an overview of the datasets included in our benchmark, which comprises a total of 17 datasets from 9 different domains. These datasets include 6 from one source, 10 from two sources and 1 from three sources. For the datasets sourced from two sources, we include 5 homogeneous and 5 heterogeneous datasets. The number of attributes per entity record ranges from 1 to 31, with an average number of duplicates per record ranging from 0.02 to 1. Different “notebook” and “altosight” datasets originate from different years of the SIGMOD Programming Contest. After preprocessing and labeling, most datasets are found to be homogeneous, especially those from DeepMatcher (Mudgal et al., 2018). For the heterogeneous datasets, we include both artificially heterogeneous datasets, constructed by renaming and merging attributes as (Wang et al., 2021b), and naturally heterogeneous datasets where raw data was available. For example, we constructed “fodors-zagats_heter” and “abt-buy_heter” by renaming and merging attributes. In contrast, “walmart-amazon_heter” comprises the original heterogeneous records with more than 15 attributes and labels from “walmart-amazon_homo”. The multi-source “movies” dataset is a merge of three datasets (IMDB-TMDB, IMDB-TVDB, and TMDB-TVDB) obtained from JedAIToolkit (Papadakis et al., 2020). Details of the datasets and processing can be found in the supplementary material.
@@ 5. Experiments
DeepBlocker Sudowoodo STransformer UniBlocker
@@
In the following section, we conduct thorough experiments to determine the effectiveness and universality of UniBlocker, as well as to identify its strengths and weaknesses in comparison to sparse blocking methods. We first evaluate the performance of UniBlocker against both domain-specific and unified sentence-based dense blocking methods. Then we compare UniBlocker with sparse blocking methods. Finally, we conduct ablation studies.
@@ 5.1. Experimental Setup
We implemented UniBlocker using Python 3.10, Pytorch 1.13, and the Transformers (Wolf et al., 2020) library. We chose to use the all-mpnet-base-v2 from STransformer (Reimers and Gurevych, 2019) as the initial pre-training weight and set the maximum sequence length to 256. We set the probability of applying literal paraphrasing actions to 0.01, the dropout probability for feature paraphrasing to 0.15, and the similarity threshold $s$ for duplicate detection to 0.85. The scaling factor $\gamma$ and margin $m$ of the Circle Loss are the default 80 and 0.4, respectively. We used the AdamW optimizer with a learning rate of 1e-5, a linear learning rate schedule without warm-up steps, and a batch size of 128 for pre-training. The total number of pre-training steps is 10,000.
@@ 5.2. Comparison with Dense Blocking Methods
UniBlocker vs Domain-Specific Methods.
To assess the effectiveness and universality of UniBlocker, we first compare UniBlocker with SOTA domain-specific self-supervised dense blocking methods. Experimental results presented in Table 3 indicate that UniBlocker outperforms previous domain-specific dense blocking methods by almost 10% mAP on average. In contrast to the lack of a clear winner between DeepBlocker and Sudowoodo, UniBlocker consistently delivers the best performance among them on almost all datasets, further proving its effectiveness by reducing the number of nearest neighbors required. These results indicate that UniBlocker not only addresses the problem of limited downstream data, but also eliminates the need for domain-specific model training. Pre-training allows deep models to reach their full potential and learn a more general representation of the structured records. In conclusion, our experimental results provide compelling evidence for the advantage of self-supervised dense blocking learning on large-scale open-domain data.
Finding 1. UniBlocker outperforms SOTA domain-specific dense blockers, showing the feasibility of developing universal dense blockers via a unified domain-independent pre-training.
@@
Blocking Workflows Sparkly3 UniBlocker UniBlocker + Sparkly3
@@
UniBlocker vs Unified Sentence-Based Models.
To explore the necessity of pre-training on structured data, we compare UniBlocker with the sentence embedding language model. The performance of STransformer, which served as our initialized model for pre-training, is detailed in Table 3. Our experimental results show that UniBlocker outperforms STransformer by over 10% mAP on average. STransformer performs relatively poorly on some less challenging but more structured datasets, such as “fodors-zagats_homo”, “abt-buy_homo”, and “walmart-amazon_homo”. This highlights the barriers to applying language models directly to blocking due to the divergences between unstructured natural language and structured records. Overall, our results emphasize the necessity of training the universal blocker on structured and semi-structured records.
Finding 2. UniBlocker outperforms sentence-based models, supporting the need to align unstructured language models with structured and semi-structured records by pre-training.
Domain-Specific Fine-tuning for UniBlocker.
To investigate the potential benefits of domain-specific fine-tuning for UniBlocker, we further fine-tune UniBlocker in the one-model-per-dataset paradigm. The mAP changes of the fine-tuned models are also presented in Table 3. Comparing the experimental results with and without fine-tuning, we observe that UniBlocker derives limited benefit from fine-tuning except for “imdb-dbpedia”. This finding suggests that the domain-independent pre-training of UniBlocker has resulted in a universal blocking capability that is sufficient to address domain-specific blocking challenges without the need for further fine-tuning. Consequently, this finding reaffirms the effectiveness of UniBlocker in solving blocking problems across multiple domains.
Finding 3. UniBlocker benefits little from domain-specific self-supervised fine-tuning, indicating that UniBlocker can learn sufficient knowledge from domain-independent pre-training.